OpenStudy (anonymous):

sin-¹(-1/2) the answer is -π/6 how do I get that answer?

OpenStudy (ash2326):

We have \[\sin^{-1} (-1/2)\]? do you know the range of \(\sin^{-1} x\)?

OpenStudy (ash2326):

Are you here @SandyHearts ?

OpenStudy (anonymous):

All it says is find the exact value of each expression. No range or anything.

OpenStudy (ash2326):

I know that but do you know about \(\sin^{-1} x\)?

OpenStudy (anonymous):

no i do not

OpenStudy (ash2326):

Do you want me to tell you about it? then the problem will be solved in seconds!!

OpenStudy (anonymous):

yes, of course.

OpenStudy (anonymous):

oh wait the anser is -π/6

OpenStudy (anonymous):

I just don't know how to get it.

OpenStudy (ash2326):

Ok, I'll explain you. You know sin x?

OpenStudy (anonymous):

what about it?

OpenStudy (ash2326):

What is the range of sin x? I mean what's its minimum value and what is its maximum value?

OpenStudy (anonymous):

0-180? i have no idea

OpenStudy (ash2326):

You know when \[x =0\longrightarrow \sin x =0\] \[x =\pi/2\longrightarrow \sin x =1\] \[x =-\pi/2\longrightarrow \sin x =-1\] Do you get this?

OpenStudy (anonymous):

yes I get this. Like in the unit circle.

OpenStudy (ash2326):

Yeah, so sin x max value is 1 and min value is -1

OpenStudy (ash2326):

You know the inverse of sin x is \(\sin^{-1}x\) Just like division is inverse of multiplication We multiply x by 2 and divide it by 2, we'll get back x Do you get this?

OpenStudy (ash2326):

You know sin x repeats its value after \(2\pi\)? So we have fixed the range and domain of \(\sin^{-1} x\) \[ \text{x can only be between -1 and 1}\] \[\text{and its range is from }\frac{-\pi}{2} \text {to}\frac{\pi}{2}\] Do you get this?

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

Ok tell me at what value of x , sin x =1/2?

OpenStudy (anonymous):

sin-¹(1/2)=x @ash2326

OpenStudy (ash2326):

I mean When \(x=60\ \sin x=\frac{\sqrt 3}{2}\) so for what value of x, sin x=1/2, check your unit circle and tell

OpenStudy (anonymous):

π/6 ?

OpenStudy (ash2326):

Yeah:D you're right so tell me know \[\sin^{-1} (-\frac{1}{2})=?\]

OpenStudy (anonymous):

-π/6

OpenStudy (ash2326):

Yay:D did you understand it?

OpenStudy (anonymous):

Yes, thank you so much!

OpenStudy (ash2326):

Welcome:D